Sisters Corp. expects to earn $6 per share next year. The firms ROE is 15% and its plowback ratio is 60%. The firms market
Answer #1

1.
=E1*(1-plowback ratio)/(market capitalization rate-ROE*plowback ratio)
=6*(1-60%)/(10%-15%*60%)
=$240.00

2.

=E1.market capitalization rate
=6/10%
=60.00

3.
=240.00-60.00
=180.00